OpenStudy (genericnoodles):

I really need assistance, any help is appreciated :) Arrange these numbers from least to greatest: 5/6, 7/9, -4/5 (These are rational numbers, and those are fractions) A.-4/5, 5/6, 7/9 B. 7/9, 5/6, -4/5 C. 5/6, 7/9, -4/5 D.-4/5, 7/9, 5/6

OpenStudy (anonymous):

Well narrow your answers down to two

OpenStudy (anonymous):

Can you?

OpenStudy (anonymous):

If not just say so

OpenStudy (genericnoodles):

Thanks for responding! I could try my best, maybe you could tell me if my method is correct?

OpenStudy (anonymous):

Of course

OpenStudy (genericnoodles):

I multiply 5 and 9, and repeat the process with the other fractions until I find the least greatest and greatest. Im having trouble because I got a similar question wrong using the same process.

OpenStudy (anonymous):

No, Okay this is how you'd want to go about. One of your numbers is a negative, right?

OpenStudy (genericnoodles):

Yes, -4/5

OpenStudy (anonymous):

Okay the other two are positive? then that narrows it down to A & D because a negative is always less than a positive

OpenStudy (genericnoodles):

The method I was using is called the Comparison property, I ment to mention it sorry!

OpenStudy (genericnoodles):

And I didn't make sense of the negative, that may have been where I messed up!

OpenStudy (anonymous):

No no just follow with me. You get that the negative is smaller than the other two?

OpenStudy (genericnoodles):

Yes, I see that.

OpenStudy (anonymous):

Okay do you know how to find the decimal of those fractions?

OpenStudy (anonymous):

the other two.

OpenStudy (genericnoodles):

I suppose so, should I make them into decimals and put them in order?

OpenStudy (anonymous):

Yes. Put them in this discussion before finalizing

OpenStudy (anonymous):

If you don't know how to just ask.

OpenStudy (genericnoodles):

Will do! One second :)

OpenStudy (genericnoodles):

I messed up Im sorry, 5/6= 0.833333 -4/5= .8 7/9= 0.777778

OpenStudy (genericnoodles):

Is this correct?

OpenStudy (anonymous):

The negative stays a negative. but you are correct. It'd be D

OpenStudy (genericnoodles):

Thank you a lot man, you have been an awesome help!

OpenStudy (anonymous):

Anytime, If you wouldn't mind 'best response' if I helped(:

OpenStudy (genericnoodles):

I'm not going to close the question just yet, can you tell me why using the comparison property wasn't very good?

OpenStudy (anonymous):

You can tag me in your questions and if im not busy doing school myself id me more than happy to try and assist.

OpenStudy (anonymous):

I'm not familiar with the term

OpenStudy (genericnoodles):

You take 2 fractions and multiply the numerator of one, and the denominator of the other, you do this with both fractions, the smaller number is the least greatest

OpenStudy (genericnoodles):

It was one of the ways taught in the lesson, I believe you can divide as well

OpenStudy (anonymous):

No clue, Thats always how I have done it. Convert it to something that is similar to them all.

OpenStudy (genericnoodles):

Well thanks for the help dude, this was killing me xD

OpenStudy (anonymous):

Haha no problem, at least you understand what to do

OpenStudy (anonymous):

Best response please